OpenStudy (anonymous):

hockey teams cores 2 points for a win, 1 point for a tie ; they won 10 more times than they tied for a total of 65 points, how many were tied?

OpenStudy (anonymous):

65 = 2 (x +10) +x 65 = 3x + 20 45 = 3x x = 15